A stretched wire between 2 points 1.0m apart is plugged by a mass of 4g, the lowest resonance frequency is 80Hz. Calculate
i- The wavelength
ii- The tension in the wire
1
Expert's answer
2018-05-28T13:22:07-0400
i.
λ=2L=2(1.0)=2.0 m.
ii.
v=λf=√(T/μ)
T=μ(λf)^2=m/L (λf)^2=0.004/1 (2∙80)^2=100 N.